# R3R: Decentralized Multi-Agent Collision Avoidance with Infinite-Horizon Safety

### v1

@@ III-A R-Boundedness
Some trajectory $\mathcal{T}=([t_{0},\infty),x(t),u(t))$ is $R$-Bounded if
$\left\|p(t)-p(t_{0})\right\|_{2}\leq R,\;\forall t\geq t_{0}.$ (15)
@@
Consider two $R$-Bounded trajectories $([t_{1},\infty),x_{1},u_{1}),\;([t_{2},\infty),x_{2},u_{2})$, and an inter-agent collision radius of $\delta<R$. Without loss of generality, let $t_{1}\leq t_{2}$. If:
$\|p_{1}(t_{1})-p_{2}(t_{2})\|\geq 2R+\delta$ (16)
Then:
$\|p_{1}(t)-p_{2}(t)\|\geq\delta,\forall t\geq t_{2}.$ (17)
@@
Since trajectories are $R$-bounded, $p_{1}(t)\in B_{1}=\mathcal{B}(p_{1}(t_{1}),R),\;p_{2}(t)\in B_{2}=\mathcal{B}(p_{2}(t_{2}),R)$, $\forall t\geq t_{2}$. Then,
$\min_{t\geq t_{2}}\left\|p_{1}(t)-p_{2}(t)\right\|$ $\geq\operatorname{dist}(B_{1},B_{2})$
$=\left\|p_{1}(t_{1})-p_{2}(t_{2})\right\|-2R\geq\delta.$
@@
Lemma˜1 provides a powerful safety condition based on the trajectories’ anchor points. However, for an agent replanning at time $t_{2}$,, the anchor point of another agent’s trajectory (created at an earlier time $t_{1}$) is not immediately relevant. To be useful, we must re-state this condition in terms of the agents’ current positions at the moment of planning.
Lemma 2 (R3R).
Consider two $R$-Bounded trajectories $([t_{1},\infty),x_{1},u_{1}),\;([t_{2},\infty),x_{2},u_{2})$, with starting times $t_{1}<t_{2}$ for some radius $R$. If at some $t\geq t_{2}$ the agents collide
$\left\|p_{2}(t)-p_{1}(t)\right\|<\delta$ (18)
@@
$\left\|p_{2}(t_{2})-p_{1}(t_{2})\right\|<3R+\delta.$ (19)
@@
$\left\|p_{2}(t)-p_{1}(t)\right\|<\delta\Rightarrow\left\|p_{2}(t_{2})-p_{1}(t_{1})\right\|<2R+\delta.$ (20)
@@
$\begin{aligned} \left\|p_{2}(t_{2})-p_{1}(t_{2})\right\|&\leq\left\|p_{2}(t_{2})-p_{1}(t_{1})\right\|+\left\|p_{1}(t_{2})-p_{1}(t_{1})\right\|\\ &<2R+\delta+R=3R+\delta\quad\qed\end{aligned}$
Lemma˜2 provides a condition under which two $R$-Bounded trajectories can collide given the state of the system at the time agent $2$ is constructing its committed trajectory. Letting agents communicate within a ball of $3R+\delta$ allows us to rule out collisions with agents outside the replanning agent’s communication range. We will thus require that the planning radius of all agents satisfy
$R^{\text{comm}}=3R^{\text{plan}}+\delta,$ (22)
which we term the R3R constraint.
III-B Decentralized Safety Protocol
We now integrate the concept of $R$-Bounded trajectories into a decentralized safety protocol. We use gatekeeper as the building block for safety, and leverage $R$-Boundedness to enable local safety verification of candidate trajectories.
Each agent $i$ executes gatekeeper independently, at discrete iterations indexed by $k_{i}\in\mathbb{N}$, which occur at times $t_{k_{i}}\in\mathbb{R}_{\geq 0}$, where $t_{k_{i}}<t_{(k+1)_{i}}$. Iteration $0_{i}$ coincides with agent $i$ entering $\mathcal{A}$. The agent is required to have a valid candidate at this time. At later iterations $k_{i}>0$, agent $i$ performs the following:
@@ III-B Decentralized Safety Protocol
    Receives the current committed trajectories $\mathcal{T}_{k_{j},j}^{\rm com}$ of all its neighbors $j\in\mathcal{N}_{i}(t_{k_{i}})$.
@@
    Plans a nominal trajectory $\mathcal{T}_{k_{i},i}^{\rm nom}$ to its goal, attempting to avoid the committed trajectories of all its neighbors.
@@
    Construct candidate trajectories as in algorithm˜1 and defined in definition˜8, until either one is found that satisfies the validity conditions of definition˜10, or none is found.
@@
    As in algorithm˜2, if a valid candidate is found, then it is committed, otherwise the previous committed trajectory is used.
This algorithm is fully decentralized, requiring only communication between neighboring agents, and asynchronous.
Input: $x_{i},g_{i},\mathcal{N}^{com},\mathcal{S}$
Result: success (bool), $\mathcal{T}^{{\rm can},T_{S}}$
$\mathcal{T}^{nom}\leftarrow\text{PlanNominal}(x_{i},g_{i},\mathcal{N}^{com},\mathcal{S})$
for $T_{S}\in[T_{H},0]$ do
 $\mathcal{T}^{{\rm can},T_{S}}\leftarrow\text{ConstructCandidate}(\mathcal{T}^{nom},T_{S})$
 if $\text{IsValid}(\mathcal{T}^{{\rm can},T_{S}},\mathcal{N}^{com},\mathcal{S})$ then
    return True, $\mathcal{T}^{{\rm can},T_{S}}$
@@
// Trigger: Agent $i$ wants to join/replan.
success, $\mathcal{T}^{can}\leftarrow\text{AttemptReplan()}$
@@
 $\mathcal{T}_{k_{i}+1,i}^{com}\leftarrow\mathcal{T}^{can}$
 // Commit to the new valid trajectory
 if agent $i\notin\mathcal{A}(t)$ then
    $\mathcal{A}(t)\leftarrow\mathcal{A}(t)\cup\{i\}$
    // Join the network
else if agent $i\in\mathcal{A}(t)$ then
 $\mathcal{T}_{k_{i}+1,i}^{com}\leftarrow\mathcal{T}_{k_{i},i}^{com}$
 // Gatekeeper: Keep the old trajectory
@@
Replanning iterations can be triggered by an internal clock (i.e. every $T_{\text{replan}}$ seconds), or by some other internal logic (i.e. when the agent is close to the switch time of its current committed trajectory). In general, the replanning times $t_{k_{i}}$ are not synchronized between agents, and can be completely independent of each other, subject to assumption˜3, defined later. No assumptions need to be placed on the frequency of replanning to ensure safety.
The safety of this protocol hinges on the IsValid check within algorithm˜1. Validity represents the conditions a candidate trajectory must satisfy in order to be considered safe. We formally define those criteria as follows.
Definition 10 (Valid).
A candidate trajectory $\mathcal{T}^{{\rm can},T_{S}}_{k_{i},i}$ as defined in definition˜8 is valid if all the following hold:
1) The trajectory remains in the safe set during $t_{k_{i}},t_{kB}]$
$x_{k,i}^{\rm can}(t)\in\mathcal{S},\forall t\in[t_{k_{i}},t_{kB}].$ (23)
2) The trajectory reaches the backup set $\mathcal{C}_{k_{i},i}$ at $t_{k_{i},B}$
$x_{k_{i},i}(t_{k_{i},B})\in\mathcal{C}_{k_{i},i},$ (24)
where $t_{k_{i},B}=t_{k_{i}}+T_{S}+T_{B}$ is the time the trajectory enters the backup set.
3) The trajectory remains within the planning bound $R^{\text{plan}}$ for all time:
[…]
$\operatorname{dist}(\mathcal{C}_{k_{i},i},p_{i}(t_{k_{i}}))\leq$ $R^{\text{plan}}.$ (26)
4) The candidate trajectory is collision free with respect to the committed trajectories of all neighboring agents,
$\|p_{k_{i},i}^{\rm can}(t)-p_{k_{j},j}^{\rm com}(t)\|\geq\delta,\forall t\in[t_{k_{i}},\infty),\forall j\in\mathcal{N}_{i}(t_{k_{i}})$ (27)
Theorem 1 (R3R - Safety of Valid Trajectories).
Consider the system described by ˜1 at some arbitrary time $t_{k}$, with the set of active agents being $\mathcal{A}(t_{k})$. Assume all agents satisfy the R3R condition (22). If at time $t_{k}$, every agent $i\in\mathcal{A}(t_{k})$ is executing a trajectory certified as valid per definition˜10 at its respective time of creation $t_{k_{i}}<t_{k}$, then the system is guaranteed to be safe for all future time. That is, $\forall t\geq t_{k}$:
$\left\|p_{i}(t)-p_{j}(t)\right\|\geq\delta,$ $\quad\forall i,j\in\mathcal{A},i\neq j,$ (28)
$x_{i}(t)\in\mathcal{S},$ $\quad\forall i\in\mathcal{A}.$ (29)
Proof.
[Proof Strategy] The proof is conducted in two parts. First, we prove inter-agent collision avoidance by contradiction, showing that a collision would violate the trajectory’s validity conditions at the time of creation. Second, we directly prove that agents remain in the safe set by construction.
[Collision Avoidance] Seeking contradiction, assume that at some $t\geq t_{k}$, distinct agents $i,j\in\mathcal{A}$ collide, i.e., $\exists t\geq t_{k}$ such that $\left\|p_{i}(t)-p_{j}(t)\right\|<\delta$. Without loss of generality, let $t_{k_{i}}<t_{k_{j}}$. [^3] In the case of $t_{k_{i}}=t_{k_{j}}$, collision is precluded by lemma 1 and assumption 3, to be defined. Both trajectories are valid by definition˜10, and thus $R$-Bounded by $R^{\text{plan}}$. Applying lemma˜2, the condition $\left\|p_{i}(t)-p_{j}(t)\right\|<\delta$ implies that at time $t_{k_{j}}$, the agents’ positions must have satisfied $\left\|p_{i}(t_{k_{j}})-p_{j}(t_{k_{j}})\right\|<3R^{\text{plan}}+\delta$, which per the R3R condition (22) is precisely $R^{\text{comm}}$, thus $i\in\mathcal{N}_{j}(t_{k_{j}})$. This is a contradiction, since by definition˜10, agent $j$’s trajectory $p_{j}(t)$ being valid at $t_{k_{j}}$ requires that it is collision free with respect to the trajectories of all agents in $\mathcal{N}_{j}(t_{k_{j}})$, and hence collision free with respect to the trajectory $p_{i}(t)$ of agent $i\in\mathcal{N}_{j}(t_{k_{j}})$, $\forall t\geq t_{k_{j}}$. This contradicts our assumption that agents $i$ and $j$ collide at some time $t\geq t_{k_{j}}$. Therefore our assumption of a collision must be false, i.e., the agents do not collide.
[Safe Set] We now show the agent remains in the safe set. By construction, a valid trajectory satisfies $x_{i}(t)\in\mathcal{S},\forall t\in[t_{k_{i}},t_{k_{i},B}]$, and $x_{i}(t)\in\mathcal{C}_{i},\forall t\geq t_{k_{i},B}$, which by definition˜5, $\mathcal{C}_{i}\subset\mathcal{S}$ thus $x_{i}(t)\in\mathcal{S},\forall t\geq t_{k_{i}}$. ∎
Theorem˜1 provides a condition for infinite-horizon safety at a single instant, given agents all track valid trajectories. We must now show all agents in the network always track valid trajectories, thus showing the theorem is satisfied after every iteration of gatekeeper. This is the principle of forward invariance, which we achieve by formalizing how an agent "commits" to a trajectory. gatekeeper’s key function is to ensure that an agent only replaces its current safe plan with a new, provably safe plan, or else continues with the old one. This is formalized with the notion of the committed trajectory, which is the trajectory tracked by the agent, and which is constructed to satisfy the validity conditions of definition˜10 as follows.
Definition 11 (Committed Trajectory).
At the $k$th iteration, let agent $i\in\mathcal{A}(t_{k_{i}})$ be at state $x_{i}(t_{k_{i}})\in\mathcal{X}$. Let $\mathcal{T}_{k_{i},i}^{\rm can}$ be the candidate trajectory constructed at time $t_{k_{i}}$ by agent $i$, as in definition˜8. Define the set of valid switch times
$\mathcal{I}_{k_{i}}=\{T_{S}\in[0,T_{H}]:\mathcal{T}_{k_{i},i}^{\rm can}\text{ is valid.}\},$ (30)
where validity is defined by definition˜10. If $\mathcal{I}_{k_{i}}\neq\varnothing$, let $T_{S}^{*}=\max\;\mathcal{I}_{k_{i}}$, then $\mathcal{T}_{k_{i},i}^{\rm com}=\mathcal{T}_{k_{i},i}^{{\rm can},T_{S}^{*}}$. Otherwise, let $\mathcal{T}_{k_{i},i}^{\rm com}=\mathcal{T}_{k_{i}-1,i}^{\rm com}$ (i.e. continue tracking the previously committed trajectory).
Assumption 3.
No two agents which can communicate update their committed trajectories simultaneously.
Remark 2.
This assumption is necessary to ensure replanning agents have accurate information their neighbors [19]. If two neighboring agents replan simultaneously, they could each construct and validate colliding candidate trajectories, as neither would have knowledge of the other. An approach to relaxing this assumption is explored in [10], uses a check-recheck scheme to ensure agents have sufficiently up-to-date information about their neighbors’. Alternatively, a randomized backoff could be implemented, such that if two agents attempt to replan simultaneously, both wait for a short, randomized period before trying again. We acknowledge this is a significant limitation and an area of future work.
Finally, we combine these elements into a single theorem that proves the forward-invariant safety of the entire protocol. This theorem asserts that as long as agents join the network and update their trajectories according to the specified rules, every agent will possess a valid committed trajectory at all times, thereby guaranteeing perpetual system-wide safety.
Theorem 2 (Forward-Invariant Safety via Update Protocols).
Consider the system described in ˜1, initialized at time $t_{0}$ with an empty set of agents $\mathcal{A}(t_{0})=\varnothing$. If all agents satisfy the R3R condition (22), perform all actions (i.e. joining and replanning) by algorithm˜2 subject to assumption˜3, then for all $t\geq t_{0}$, the system is guaranteed to remain safe: (1). Collision Avoidance: $\left\|p_{i}(t)-p_{j}(t)\right\|\geq\delta$ for all distinct $i,j\in\mathcal{A}(t)$, (2). Safe Set: $x_{i}(t)\in\mathcal{S},\;\forall i\in\mathcal{A}(t)$.
Proof.
The proof is by induction.
[Initial Conditions] The initial condition $\mathcal{A}(t_{0})=\varnothing$ is trivially safe. Consider the first agent $i$ to join the network at time $t_{0_{i}}$. By algorithm˜2, agent $i$ can only join the network if algorithm˜1 finds a valid candidate trajectory $\mathcal{T}_{0,i}^{{\rm can},T_{S}}$. By definition˜10, this trajectory must remain in the safe set $\mathcal{S}$ for all time, and because there are no other agents, the inter-agent collision constraint is trivially satisfied.
[Inductive Hypothesis] Assume at some time $t_{k}$, all agents in the network $\mathcal{A}(t_{k})$ are tracking committed trajectories that were valid at their respective times of construction.
[Inductive Step] We must show the system remains safe after the next replanning event at time $t_{k+1}>t_{k}$. This can either be a new agent joining, or existing agent replanning.
Case 1: A New Agent Joins Agent $j\notin\mathcal{A}(t_{k+1})$ must execute algorithm˜2, and construct a valid committed trajectory $\mathcal{T}_{0,j}^{\rm com}$. By theorem˜1, because agent $j$’s trajectory and all existing agents trajectories are valid, the entire system is guaranteed to be safe.
Case 2: Existing Agent Replans Assume agent $i\in\mathcal{A}(t_{k+1})$ executes algorithm˜2. If the AttemptReplan algorithm fails, the agent continues tracking its previous committed trajectory. The set of active trajectories in the system is unchanged, so the system remains safe by the inductive hypothesis.
If AttemptReplan succeeds,the agent adopts a new Valid committed trajectory, $\mathcal{T}_{k+1,i}^{\rm com}$, which as before is guaranteed to be safe by the validity conditions as shown in theorem˜1.
Thus, since the system starts in a safe state, and the protocols ensure state transitions are required to be safe, the system must remain safe for all future time. ∎
IV Simulation Results

### v2

@@ III-B $R$-Boundedness
A trajectory $\mathcal{T}=([t_{0},\infty),{\rm x}(t),{\rm u}(t))$ is R-Bounded if:
$\left\|{\rm p}(t)-{\rm p}(t_{0})\right\|_{2}\leq R,\;\forall t\geq t_{0}.$ (15)
@@
Consider two $R$-Bounded trajectories $([t_{1},\infty),{\rm x}_{1},{\rm u}_{1}),\;([t_{2},\infty),{\rm x}_{2},{\rm u}_{2})$, and suppose the system has an inter-agent collision radius of $\delta<R$. Without loss of generality, let $t_{1}\leq t_{2}$. If
$\left\|{\rm p}_{1}(t_{1})-{\rm p}_{2}(t_{2})\right\|_{2}\geq 2R+\delta,$ (16)
then
$\left\|{\rm p}_{1}(t)-{\rm p}_{2}(t)\right\|_{2}\geq\delta,\forall t\geq t_{2}.$ (17)
@@
Since trajectories are $R$-Bounded, ${\rm p}_{1}(t)\in B_{1}=\mathcal{B}_{R}({\rm p}_{1}(t_{1})),\;{\rm p}_{2}(t)\in B_{2}=\mathcal{B}_{R}({\rm p}_{2}(t_{2}))$, $\forall t\geq t_{2}$. Then,
$\min_{t\geq t_{2}}\left\|{\rm p}_{1}(t)-{\rm p}_{2}(t)\right\|_{2}$ $\geq\min_{p_{1}\in B_{1},\;p_{2}\in B_{2}}\left\|p_{1}-p_{2}\right\|_{2}$
$=\left\|{\rm p}_{1}(t_{1})-{\rm p}_{2}(t_{2})\right\|_{2}-2R\geq\delta.$
@@
While lemma˜1 establishes safety based on the trajectory anchor points, we must translate this into a condition based on the agents’ current positions at the moment of replanning.
Lemma 2 (Sufficient Communication Radius).
Consider two $R$-Bounded trajectories $([t_{1},\infty),{\rm x}_{1},{\rm u}_{1}),\;([t_{2},\infty),{\rm x}_{2},{\rm u}_{2})$, with starting times $t_{1}<t_{2}$ for some radius $R$. If at some $t\geq t_{2}$ the agents collide,
$\left\|{\rm p}_{2}(t)-{\rm p}_{1}(t)\right\|_{2}<\delta,$ (18)
@@
$\left\|{\rm p}_{2}(t_{2})-{\rm p}_{1}(t_{2})\right\|_{2}<3R+\delta.$ (19)
Thus, if trajectories are bounded by a radius $R^{\text{plan}}$ satisfying
$R^{\text{comm}}=3R^{\text{plan}}+\delta,$ (20)
which we term the R3R condition, then agents $1$ and $2$ must be neighbors at $t_{2}$, i.e., $(1,2)\in E(t_{2})$.
@@
$\left\|{\rm p}_{2}(t)-{\rm p}_{1}(t)\right\|_{2}<\delta\Rightarrow\left\|{\rm p}_{2}(t_{2})-{\rm p}_{1}(t_{1})\right\|_{2}<2R+\delta.$ (21)
@@
$\left\|{\rm p}_{2}(t_{2})-{\rm p}_{1}(t_{2})\right\|_{2}$ $\leq\left\|{\rm p}_{2}(t_{2})-{\rm p}_{1}(t_{1})\right\|_{2}+\left\|{\rm p}_{1}(t_{2})-{\rm p}_{1}(t_{1})\right\|_{2}$
$<2R+\delta+R=3R+\delta.$
By letting $R=R^{\text{plan}}$ and applying the condition $R^{\text{comm}}=3R^{\text{plan}}+\delta$, we have
$\left\|{\rm p}_{2}(t_{2})-{\rm p}_{1}(t_{2})\right\|_{2}$ $\leq 3R^{\text{plan}}+\delta,$
$=R^{\text{comm}}.$
And therefore, by Def.˜3, $(1,2)\in E(t_{2})$. ∎
This extra $R$ relative to lemma˜1 arises from the fact that agent $1$ may move up to a distance $R$ between $t_{1}$ and $t_{2}$, and results in Eq. (20), the R3R condition. This informs a constraint on trajectories which, if satisfied by all agents in the system, enables global safety certification with local information.
III-C Decentralized Safety Protocol
We now integrate this constraint with gatekeeper to establish a formally safe decentralized protocol. We use gatekeeper as the building block for safety, and leverage $R$-Boundedness to enable local safety verification of candidate trajectories.
When an agent $i$ joins the network, which we term iteration $0_{i}$, the agent is required to have a valid candidate. At later iterations $k_{i}>0$, agent $i$ will:
@@ III-C Decentralized Safety Protocol
    Receive the current committed trajectories $\mathcal{T}_{k_{j},j}^{\rm com}$ of all neighbors $j\in\mathcal{N}_{i}(t_{k_{i}})$.
@@
    Plan a nominal trajectory $\mathcal{T}_{k_{i},i}^{\rm nom}$, attempting to avoid static obstacles and the committed trajectories of neighbors.
@@
    Construct candidates using algorithm˜1 until one is found which satisfies the validity conditions of Def. 9, or the algorithm returns.
@@
    If a valid candidate is found, commit the trajectory, otherwise continue along $\mathcal{T}_{k_{i}-1,i}^{\rm com}$.
Remark 1.
No conditions are placed on the frequency of replanning iterations which can be triggered independently and asynchronously by agents based on internal logic.
We now formalize the conditions under which a candidate trajectory can be certified as safe.
Definition 9 (Valid Candidate Trajectory).
At time $t_{k_{i}}$, let agent $i$ hold a previously committed trajectory $\mathcal{T}^{{\rm com}}_{k_{i}-1,i}$. A candidate trajectory $\mathcal{T}^{{\rm can},\Delta t_{S}}_{k_{i},i}$ is valid if it satisfies:
  1. 1.
    Safe Set: ${\rm p}_{k_{i},i}^{{\rm can},\Delta t_{S}}(t)\in\mathcal{S},\forall t\in[t_{k_{i}},t_{k_{i},B}]$.
  2. 2.
    Backup Reachability:
    ${\rm x}_{k_{i},i}^{{\rm can},\Delta t_{S}}(t_{k_{i},B})\in\mathcal{C}_{k_{i},i},$ (23)
    where $t_{k_{i},B}=t_{k_{i}}+\Delta t_{S}+\Delta t_{B}$.
  3. 3.
    Backup Separation:
    $\inf_{\begin{subarray}{c}x_{1}\in\mathcal{C}_{k_{i},i}\\ x_{2}\in\mathcal{C}_{k_{j},j}\end{subarray}}\left\|\Pi(x_{1})-\Pi(x_{2})\right\|_{2}\geq\delta,\;\forall j\in\mathcal{N}_{i}(t_{k_{i}}).$ (24)
  4. 4.
    R-Boundedness:
    $\left\|{\rm p}_{k_{i},i}^{{\rm can},\Delta t_{S}}(t)-{\rm p}_{i}(t_{k_{i}})\right\|_{2}$ $\leq R^{\text{plan}},\forall t\in[t_{k_{i}},t_{k_{i},B}],$
    $\sup_{x\in\mathcal{C}_{k_{i},i}}\left\|\Pi(x)-{\rm p}_{i}(t_{k_{i}})\right\|_{2}$ $\leq R^{\text{plan}}.$
  5. 5.
    Collision Avoidance: $\forall t\in[t_{k_{i}},\max\{t_{k_{i},B},t_{k_{j},B}\}]$,
    $\left\|{\rm p}_{k_{i},i}^{{\rm can},\Delta t_{S}}(t)-{\rm p}_{k_{j},j}^{\rm com}(t)\right\|_{2}\geq\delta,\;\forall j\in\mathcal{N}_{i}(t_{k_{i}}).$ (25)
Remark 2.
The conditions given by (24) and (25) together provide an infinite-horizon safety guarantee. Condition (25) ensures finite-horizon collision avoidance while the agents are actively traversing their trajectories, and condition (24) ensures that once both agents reach their respective backup sets, they will remain permanently separated for all future time.
Theorem 1 (R3R - Safety of Valid Trajectories).
If all agents $i\in\mathcal{A}(t_{k})$ choose $R^{\text{plan}}$ satisfying the R3R condition (20), and execute trajectories $\mathcal{T}_{k_{i},i}^{\rm com}$ that were valid by Def. 9 when committed at times $t_{k_{i}}<t_{k}$, then $\forall t\geq t_{k}$, agents executing these committed trajectories remain safe:
$\left\|{\rm p}_{i}(t)-{\rm p}_{j}(t)\right\|_{2}\geq\delta,$ $\quad\forall i,j\in\mathcal{A},i\neq j,$ (26)
${\rm p}_{i}(t)\in\mathcal{S},$ $\quad\forall i\in\mathcal{A}.$ (27)
Proof.
Collision avoidance. Seeking contradiction, assume that at some $t\geq t_{k}$, distinct agents $i,j\in\mathcal{A}$ collide, i.e., $\exists t\geq t_{k}$ such that $\left\|{\rm p}_{i}(t)-{\rm p}_{j}(t)\right\|_{2}<\delta$. Without loss of generality, let $t_{k_{i}}<t_{k_{j}}$. Both trajectories are valid by Def.˜9, and thus $R$-Bounded by $R^{\text{plan}}$. Applying lemma˜2, the condition $\left\|{\rm p}_{i}(t)-{\rm p}_{j}(t)\right\|_{2}<\delta$ implies that at time $t_{k_{j}}$, the agents’ positions must have satisfied $\left\|{\rm p}_{i}(t_{k_{j}})-{\rm p}_{j}(t_{k_{j}})\right\|_{2}<3R^{\text{plan}}+\delta$, which per the R3R condition (20) is precisely $R^{\text{comm}}$, thus $i\in\mathcal{N}_{j}(t_{k_{j}})$. This is a contradiction, since by Def.˜9, agent $j$’s trajectory ${\rm p}_{j}(t)$ being valid at $t_{k_{j}}$ requires that it is collision-free with respect to the trajectories of all agents in $\mathcal{N}_{j}(t_{k_{j}})$, and hence collision-free with respect to the trajectory ${\rm p}_{i}(t)$ of agent $i\in\mathcal{N}_{j}(t_{k_{j}})$, $\forall t\geq t_{k_{j}}$.
Safe set. A valid trajectory satisfies ${\rm p}_{i}(t)\in\mathcal{S},\forall t\in[t_{k_{i}},t_{k_{i},B}]$, and ${\rm x}_{i}(t)\in\mathcal{C}_{i},\forall t\geq t_{k_{i},B}$, which by Def.˜5, implies $\Pi(\mathcal{C}_{i})\subset\mathcal{S}$, and thus ${\rm p}_{i}(t)\in\mathcal{S},\forall t\geq t_{k_{i}}$. ∎
Theorem˜1 shows that the conditions outlined in Def. 9, which rely only on local information, still provide guarantees with respect to the rest of the agents in the network. However, this theorem assumes the committed trajectories of each agent are held constant, and thus remain $R$-Bounded around the same anchor point. In order to make progress, agents must continually update their committed trajectories. We now introduce the necessary machinery to ensure all agents in the network always track valid trajectories, guaranteeing the conditions of theorem˜1 are satisfied after every iteration of gatekeeper.
Definition 10 (Committed Trajectory).
At the $k_{i}$-th iteration, let agent $i\in\mathcal{A}(t_{k_{i}})$ be at state ${\rm x}_{i}(t_{k_{i}})\in\mathcal{X}$, and $\mathcal{T}_{k_{i},i}^{\rm can}$ be the candidate trajectory. Define the set of valid switch times:
$\mathcal{I}_{k_{i}}=\{\Delta t_{S}\in[0,\Delta t_{H}]:\mathcal{T}_{k_{i},i}^{\rm can}\text{ is valid.}\},$ (28)
where validity is defined by Def.˜9. If $\mathcal{I}_{k_{i}}\neq\varnothing$, let $\Delta t_{S}^{*}=\max\;\mathcal{I}_{k_{i}}$, then $\mathcal{T}_{k_{i},i}^{\rm com}=\mathcal{T}_{k_{i},i}^{{\rm can},\Delta t_{S}^{*}}$. Otherwise, let $\mathcal{T}_{k_{i},i}^{\rm com}=\mathcal{T}_{k_{i}-1,i}^{\rm com}$ (i.e. continue tracking the previously committed trajectory).
Assumption 3.
No two agents which can communicate update their committed trajectories simultaneously, i.e.,
$t_{k}=t_{k_{i}}=t_{k_{j}}\implies(i,j)\notin E(t_{k}).$
Remark 3.
This assumption is necessary to ensure replanning agents have accurate information about their neighbors [19], preventing agents from simultaneously constructing mutually invalidating candidates. Under the conditions of assumption˜2, this is a mild assumption, however we acknowledge handling the effects of non-ideal network communication as a direction for future work. An approach to relaxing this assumption is explored in [10], using a check-recheck scheme to ensure agents have sufficiently up-to-date information about their neighbors.
The algorithmic steps of checking validity by Def.˜9, and constructing committed trajectories are shown in algorithms˜1 and 2.
Input: $x_{i},g_{i},\mathcal{N}^{{\rm com}},\mathcal{S}$
Result: success (bool), $\mathcal{T}^{{\rm can},\Delta t_{S}}$
$\mathcal{T}^{{\rm nom}}\leftarrow\text{PlanNominal}(x_{i},g_{i},\mathcal{N}^{{\rm com}},\mathcal{S})$
for $\Delta t_{S}=\Delta t_{H},\;\Delta t_{H}-\Delta t,\dots,0$ do
 $\mathcal{T}^{{\rm can},\Delta t_{S}}\leftarrow\text{ConstructCandidate}(\mathcal{T}^{{\rm nom}},\Delta t_{S})$
 if $\text{IsValid}(\mathcal{T}^{{\rm can},\Delta t_{S}},\mathcal{N}^{{\rm com}},\mathcal{S})$ then
    return True, $\mathcal{T}^{{\rm can},\Delta t_{S}}$
@@
Remark 4.
At each iteration, when constructing candidate trajectories, agents must choose some backup set $\mathcal{C}_{k_{i}}$. The same backup set can be used for each candidate, multiple can be evaluated, or its location can be optimized as part of the planning process. Agents include and transmit a representation of this backup set alongside their committed trajectories.
// Trigger: Agent $i$ join/replan.
// Receive neighbor trajectories.
$\mathcal{N}^{\rm com}\leftarrow\{\mathcal{T}_{k_{j},j}^{\rm com}\}_{j\in\mathcal{N}_{i}(t_{k_{i}+1})}$
success, $\mathcal{T}^{{\rm can}}\leftarrow\text{AttemptReplan}(x_{i},g_{i},\mathcal{N}^{\rm com},\mathcal{S})$
@@
 $\mathcal{T}_{k_{i}+1,i}^{{\rm com}}\leftarrow\mathcal{T}^{{\rm can}}$ // Commit trajectory.
 if $i\notin\mathcal{A}(t_{k_{i}+1})$ then
    $\mathcal{A}(t_{k_{i}+1})\leftarrow\mathcal{A}(t_{k_{i}+1})\cup\{i\}$ // Join.
[…]
else if $i\in\mathcal{A}(t_{k_{i}+1})$ then
 $\mathcal{T}_{k_{i}+1,i}^{{\rm com}}\leftarrow\mathcal{T}_{k_{i},i}^{{\rm com}}$ // Keep old trajectory
@@
Finally, we combine these elements into a single theorem that proves the forward-invariant safety of the entire system. This theorem asserts that as long as agents join the network and update their trajectories according to the specified rules, every agent will possess a valid committed trajectory at all times, thereby guaranteeing perpetual system-wide safety.
Theorem 2 (Forward-Invariant Safety via Update Protocols).
Assume $\mathcal{A}(t_{0})=\varnothing$. If all agents satisfy (20), and join and replan by algorithm˜2, then for all $t\geq t_{0}$, the system remains safe:
$\left\|{\rm p}_{i}(t)-{\rm p}_{j}(t)\right\|_{2}\geq\delta,\quad$ $\forall i\neq j\in\mathcal{A}(t),$ (29)
${\rm p}_{i}(t)\in\mathcal{S},\quad$ $\forall i\in\mathcal{A}(t).$ (30)
Proof.
We prove safety by induction over update events (joining or replanning) generated by algorithm˜2.
Base case. At time $t_{0}$, the active set is empty, $\mathcal{A}(t_{0})=\varnothing$, hence the safety conditions hold trivially. When the first agent $i$ joins at time $t_{0_{i}}$, algorithm˜2 requires it to commit a trajectory $\mathcal{T}_{0_{i},i}^{{\rm com}}$ satisfying Def.˜9. In particular, ${\rm p}_{i}(t)\in\mathcal{S}$ for all $t\geq t_{0_{i}}$, and collision avoidance is vacuous since no other agents are present.
Inductive hypothesis. Assume that at some update time $t_{k}$, every agent $i\in\mathcal{A}(t_{k})$ is executing a committed trajectory $\mathcal{T}_{k_{i},i}^{{\rm com}}$ that was valid at $t_{k_{i}}$. Then, by theorem˜1, the system is safe for all $t\geq t_{k}$.
Inductive step. Consider the next event at time $t_{k+1}>t_{k}$.
Case 1: Agent joining. If a new agent $i\notin\mathcal{A}(t_{k})$ joins at $t_{k+1}$, then by algorithm˜2 it may enter only after committing a valid trajectory $\mathcal{T}_{0_{i},i}^{{\rm com}}$ with respect to the currently committed trajectories of all $j\in\mathcal{N}_{i}(t_{k+1})$, and safety follows from theorem˜1.
Case 2: Agent replanning. If an agent $i\in\mathcal{A}(t_{k})$ replans at $t_{k+1}$, then either:
(i) AttemptReplan fails, in which case agent $i$ continues executing its previous committed trajectory, and the system remains safe by the inductive hypothesis; or
(ii) AttemptReplan succeeds, in which case agent $i$ commits a new trajectory $\mathcal{T}_{k_{i}+1,i}^{{\rm com}}$ that satisfies Def.˜9. Thus, safety again follows from theorem˜1.
Thus, safety is preserved across every update event. Since $t_{k}\to\infty$, the system remains safe for all $t\geq t_{0}$. ∎
An example of a system evolving under this protocol is given in Table 2. We see by repeatedly replanning, agents are able to shift their anchor points and certify trajectories that enable them to safely progress towards their goal states.
@@
If an agent elects to replan at times $t_{k_{i}+1}<t_{k_{i}}+\Delta t_{S}$ at each iteration, and is able to successfully construct a valid committed trajectory each time, it will never execute its backup controller. Therefore, an agent could avoid ever deviating from a nominal, goal-oriented trajectory, unless necessary to guarantee safety.
IV Simulation Results and Analysis